A class that defines an inner class iterator should implement the

Group of answer choices

Iteration interface

none of them

ListIterator interface

Iterator interface

When an ADT has too many operations, it is commonly known as

Group of answer choices

interface exhaustion

interface bloat

operation overload

method overload

What is the value of the following postfix expression: a b - c d / +

Assume a = 10, b = 6, c = 12, and d = 2

Group of answer choices

6

12

10

cannot be determined

What is the value of the following postfix expression: a b - c - d / +

Assume a = 10, b = 6, c = 12, and d = 2

Group of answer choices

cannot be determined

10

12

6
Answer #1

1.)

  • An inner class iterator must implement the Iterator interface
  • Answer: Iterator interface

2.)

  • When an abstract data type has more number of operations on a single data, it is called interface bloat.
  • Answer: Interface bloat

3.)

  • a b - c d / + can be conveted to infix notation as (a-b)+(c/d)
  • (a-b)+(c/d) = (10-6) + (12/2) = 4 + 6 = 10
  • Answer: 10

4.)

  • a b - c - d / + is not a valid expression, it does not have enough number of operands to perform the given operation
  • Answer: cannot be determined